OpenStudy (anonymous):

I just wanna see if I'm doing this rite. (7x -4)(7x+7) 28x 49x Did I do that correct?

OpenStudy (anonymous):

49x^2+49x-28x-28

OpenStudy (texaschic101):

(7x - 4)(7x + 7) 49x^2 + 49x - 28x - 28 49x^2 + 21x - 28

OpenStudy (anonymous):

How did you get the 2

OpenStudy (anonymous):

texaschic101

OpenStudy (texaschic101):

what 2 are you talking about ? Oh...you mean the 2 in 49x^2 ? well...7x * 7x....you not only multiply the 7's, you have to multiply the x's also. So 7x * 7x = 49x^2. Understand ?

OpenStudy (texaschic101):

you see, if you just have 7x * 7, that equals 49x, but if you have 7x * 7x, that equals 49x^2

OpenStudy (anonymous):

Yes ok so did you multiply 7x +7x+7x to get the 21x -28

OpenStudy (texaschic101):

are you familiar with the FOIL method ?

OpenStudy (anonymous):

No I'm just learning all this

OpenStudy (texaschic101):

FOIL stands for first, outside, inside, last (7x - 4)(7x + 7) First -- multiply first terms in each set...7x * 7x = 49x^2 Outside -- multiply outside terms in each set ...7x * 7 = 49x Inside -- multiply inside terms in each set...-4 * 7x = -28x Last -- multiply last terms in each set ...-4 * 7 = -28 now put it all together... 49x^2 + 49x - 28x - 28 -- combine like terms 49x^2 + 21x - 28 does that make it any clearer ?

OpenStudy (anonymous):

Yes I'm going to rite this down. Thank you. Do you mind if I call on you again?